OpenStudy (anonymous):

what is the prime factorization of 2500 using exponents

OpenStudy (kc_kennylau):

\[2500\]\[=25\times100\] Can you continue? :)

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

well, I appreciate It but I need exponents Thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks sooo Much !!! :)

OpenStudy (isaiah.feynman):

You can call me GOD.

OpenStudy (kc_kennylau):

it's actually not correct, since 10 is not a prime

OpenStudy (kc_kennylau):

\[2500\]\(=25\times100\) (extract the zeros)\[=5^2\times10^2\]\[=5^2\times2^2\times5^2\]\[=2^2\times5^4\] Learn more: https://www.youtube.com/watch?v=ZKKDTfHcsG0